Find all solutions to the following linear congruences. (15 points)

(a) 2x ≡ 5 (mod 7).

(b) 6x ≡ 5 (mod 8).

(c) 19x ≡ 30 (mod 40).

Show all the steps taken in neat English to receive a positive review
Answer #1

Solution:-

ca 2x = 5 (mod 7 I Solution eusts (2,7) 75 By solving the B Diophantine Equation 2 x - 7 :-7y=5 By Eue le dian algorithm 7 =by 6 x Ē 5 (modo) By Extended Equ nation a Eucleaian rithm for bmod in) have algoer De z we to + mq gea (am 6 pt 89 Pt 89 ged(С) 19x = 30 (под 10 ) ( 17, Чо) = 1 | 30 I so lutnon exists = Зо І Х Diophantine Equation: ГЧx И На By Euclecian algorithm